Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 128,000 $ 135,000
Accounts receivable 341,000 470,000
Inventory 571,000 486,000
Plant and equipment, net 850,000 839,000
Investment in Buisson, S.A. 399,000 430,000
Land (undeveloped) 250,000 252,000
Total assets $ 2,539,000 $ 2,612,000
Liabilities and Stockholders' Equity
Accounts payable $ 389,000 $ 350,000
Long-term debt 1,041,000 1,041,000
Stockholders' equity 1,109,000 1,221,000
Total liabilities and stockholders' equity $ 2,539,000 $ 2,612,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,966,000
Operating expenses 4,072,120
Net operating income 893,880
Interest and taxes:
Interest expense $ 119,000
Tax expense 198,000 317,000
Net income $ 576,880


The company paid dividends of $464,880 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Do not round intermediate calculations and round your final answers to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,539,000-250,000-399,000)+(2,612,000-252,000-430,000]/2

= 1,910,000

Margin = Net operating income / sales

= 893,880/4,966,000

= 18%

Turnover = Sales / Average operating assets

= 4,966,000/1,910,000

= 2.6

ROI = margin *turnover

= 18*2.6

= 46.80%

Residual income = Net operating income - required return

= 893,880 - (1,910,000*15%)

= 607,380

Comment if you face any issues